3. A review of voter registration records in a small town yielded the following table of the number of males and females registered as Democrat, Republic, or other affiliation: Democrat Mal Female Other a. Solve for x. b. What is the probability that the person is male? c.What is the probability that a randomly chosen person is female, given that they are Other affiliation? A researcher measured a stationary pulse rate per minute for a random sample of 54 adult males and found:75, s 12Using the 68-95-99 rule, between which two values do Tmost stationary pulse rates fall between, ie. which values contain 99.7% of the data? 4. 1.1 pounds. Jonathan computed a z-score,z 1.3, for his dog Lady. What is Ladys weight? Test grades in MAT 117 class have a severely left skewed distribution. A professor computed two statistics, 78 and the 70, from his test grades in one of his classes. Which of the two numbers is the mean, which one is the median? 6. Suppose the average two-year-old girl weighs 12.6 kg with standard deviation 1.4 kg. Compute the z-score for a 2 year old girl that weights 15.6 kg, round it to 2 decimal places. 7.
Answer #1

3 a) X = 1100 - 400 -400 = 300

b) probability that the person is male = 900/2000 = 0.45

c) Probability that a randomly chosen person is female, given that they are 'other affiliated = P( Female and Other)/ P(Other)

300 200030.60 00/2000.60